Dorothy owes $11,000 on two credit cards with annual interest rates 9% and 12 %. If she pays $1,140 in interestfor the year, how much does she owe on each card ?

Answer by mananth(16949)

Card I 9.00% per annum ------------- Amount invested =x
Card II 12.00% per annum ------------ Amount invested = y
11000
Interest----- 1140

Card I 9.00% per annum ---x
Card II 12.00% per annum ---y
Credit on both cards
x + 1 y= 11000 -------------1
Interestto be paid on both
9.00% x + 12.00% y= 1140
Multiply by 100
9 x + 12 y= 114000.00 --------2
Multiply (1) by -9
we get
-9 x -9 y= -99000.00
Add this to (2)
0 x 3 y= 15000
divide by 3
y = 5000
Card I 9.00% $ 6000
Card II 12.00% $ 5000

CHECK
6000 --------- 9.00% ------- 540.00
5000 ------------- 12.00% ------- 600.00
Total -------------------- 1140.00

m.ananth@hotmail.ca